Billy likes to go cycling.
His bike has wheels of diameter 75cm.
He has invented a counter for his bike,
which counts the number of revelations the wheel makes.
One day the counter shows 240 revolutions.
How far has billy cycled?
give your answer to the nearest meter.

Answer by richwmiller(17219)

you need to find the circumference of the wheel from the diameter
diameter = 2*pi
C=2*pi*r
C=pi*d
distance=240*d*pi
56549 cm
565.49 m
to the nearest meter
565 m